OpenStudy (pphalke):

what is the monthly payment at 5.50% and 5.25% Juan wishes to borrow $185,000 to purchase a condo. He will get a 20 year mortgage, and the bank is offering a rate of 5.50%. if he buys down the rate to 5.25%, how much will he reduce his monthly payment by? if he pays $1,850 for the single point, how long will it take for him to recover the cost? i got the answer of $25.98 as the reduction and 71.15 months as the time it will take to recover the cost

OpenStudy (pphalke):

@sweetburger

OpenStudy (mathmate):

Yes, 25.98 is correct. (I got 25.9798). and the months is 1850/25.9798=71.21 months, almost the same as yours. Well done!

OpenStudy (pphalke):

but what is the monthly payment for the rates i listed? 5.50 and 5.25

OpenStudy (mathmate):

You need to use the formula you got in a previous question, or use mine. They give the same results. If you have difficulty getting the right numbers, it's just you are not used to the calculator (especially the brackets). Using my formula, it is monthly payment=\(\Large \frac{PR^n*(R-1)}{R^n-1}\) Where P=185000 R=1+0.055/12 (monthly interest+1) n=number of months Try this formula and tell me what you get.

OpenStudy (mathmate):

What kind of calculator do you use?

OpenStudy (pphalke):

google, web 2.0 and a calculator site found in the book. how do i calculate monthly payment

OpenStudy (mathmate):

If you give me the exact site, I can show you what to enter, and why!

OpenStudy (pphalke):

www.interest.com/calculators

OpenStudy (mathmate):

Are you supposed to use specialized calculators, or you use a regular calculator and a formula?

OpenStudy (pphalke):

scientific I cant find any where to use to calculate monthly payment for rates

OpenStudy (mathmate):

Then you can use Google to do calculations. I'll try it out and get back to you.

OpenStudy (pphalke):

ok thanks your very kind I am stuck so I know now to come here or pa to get help I don't need other resources either, I got things I can come to get help on. simple practice quizzes I am coming here, its six week class so it wont be too bad

OpenStudy (mathmate):

If you enter 185000*(1+0.055/12)^240*(0.055/12)/((1+0.055/12)^240-1) in Google, you will get 1272.59151954 which is correct to the last place.

OpenStudy (mathmate):

You can then repeat the same for 5.25%

OpenStudy (pphalke):

I found a loan calculator it helped

OpenStudy (mathmate):

But if you take an exam, you still have to use a scientific calculator!

OpenStudy (mathmate):

The problem with loan calculators is you have to check very carefully. Put in the problem to see if it gives the same result.

OpenStudy (mathmate):

Sorry, gtg. Will continue tomorrow.